OpenStudy (anonymous):

Use seperation of variables to solve the initial value problem: dy/dx= (y+5)(x+2) and y=1 when x=0

OpenStudy (anonymous):

\[\frac{dy}{dx}=\frac{y+5}{x+2}\]\[\frac{dy}{y+5}=\frac{dx}{x+2}\]\[\int\limits_{}^{}\frac{dy}{y+5}=\int\limits_{}^{}\frac{dx}{x+2}\]Does that help?

OpenStudy (anonymous):

Its actually multiplication, not division :(

OpenStudy (anonymous):

(y+5)*(x+2)

OpenStudy (anonymous):

Ahh, my bad. So it would instead be:\[\int\limits_{}^{}\frac{dy}{y+5}=\int\limits_{}^{}(x+2)dx\]

OpenStudy (anonymous):

Yeah I got that, would it then be ln(y+5)= 1/2x^(2) + 2x +C ?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Does c = ln6?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Okay Awesome! The answer showed an equation. Do you know how I get that?

OpenStudy (anonymous):

You want an equation for y in terms of x?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

\[\ln(y+5)=\frac{x^{2}}{2}+2x+\ln(6)\]\[e^{\ln(y+5)}=e^{ \frac{x^{2}}{2}+2x+\ln(6)}\]\[y+5=e^{ \frac{x^{2}}{2}+2x} \times e^{\ln(6)}\]\[y=6 \times e^{ \frac{x^{2}}{2}+2x} -5\]

OpenStudy (anonymous):

Will you Marry me?

OpenStudy (anonymous):

I'm already engaged to someone else. :P

OpenStudy (anonymous):

You Jerk, Huge thanks though!

OpenStudy (anonymous):

You're welcome! :P